# The global communication pathways of the human brain transcend the cortical - subcortical - cerebellar division

### v1

@@ Results
Figure 4: Hierarchical centralization of cross-modular connectivity. a, Analysis of the 100 lead hub ROIs of the brain-wide network. Bar panels display their degree—sorted in decreasing order—with the bars colored according to the community to which the ROI belongs, in the three different divisions. Pie charts show the fraction of leading hubs among the subsequent communities. b, Evidence for a rich-club supraorganization among the leading hubs. Left panel, conceptual illustration of a rich-club, formed when the leading hubs are densely connected with each other. Right panel, evolution of the link density for the subnetworks formed by ROIs with degree larger than $k$ (nodes with degree smaller than $k$ are removed), for all possible thresholds of $k=0,1,\ldots,718$. Red curve is the evolution for the brain-wide network, gray curve is for degree-conserving randomized samples (100 independent realizations each point). The threshold for $k=0$ represents the entire network, thus, the curves begin from a density of 0.2. Their monotonic increase up to a density of 1.0 evidences that the leading hubs eventually form an all-to-all connected core, centralizing the communication pathways. c, Topographical location of the rich-club hubs and their mutual links, for two different density thresholds: at $k>256$ (subgraph density of 0.93) and $k>300$ (subgraph density of 0.99). Cortical ROIs are shown as green dots and subcortical ROIs in orange. d, The composition of the rich-club at different levels of internal density, from 0.8 to 1.0 (degree thresholds, $k\geq 208$ to $k=365$), displayed for the anatomical (left or right cortex, subcortex, cerebellum) and the network divisions (C1 – C5). Bars indicate the number of ROI hubs forming the rich-club at each density level, and how many of them belong to the different anatomical or network communities. e, Schematic representation of the modular and hierarchical organization of the brain-wide structural connectivity, in the light of either the anatomical or the optimal network divisions.
@@
The results of Fig. 3 highlighted the presence of hub ROIs in diverse anatomical structures. We now inquire whether these centralizing hubs form additional structures that help elucidate the global network architecture. Focusing on the distribution of the 100 most connected ROIs, Fig.4a, it shows that the leading hubs are mainly subcortical (60%), followed by cortical hubs (40%) and no presence of cerebellar ROIs among the top 100. A closer look at the subcortex (middle panel) also confirms that the subcortical hubs are very much distributed among the different subcortical structures (regardless of their size) and thus, no single subcortical structure can be considered responsible for the centralization of the communication pathways. This distribution of hub regions is more evenly observed in the light of the network-based division, with the communities C1 – C4 containing 19% – 29% of the leading hubs each.
@@
This resonates with previous observations, where cortical hubs were found to be spatially dispersed instead of forming a localized anatomical structure of their own Zamora_Hubs_2010 ; Zamora_FrontReview_2011 ; Heuvel_HubsHuman_2011 . And this, despite that cortical hubs were found to be densely interconnected, forming a communication core, also known as a rich-club supracommunity, Fig. 4b (left panel). Here too, we identify clear evidence that the brain-wide hubs form a rich-club suprastructure. By iteratively removing the ROIs with less connections, the internal density of the remaining subnetwork (made of the ROIs with larger degree) rapidly increases, Fig. 4b (right panel). There are 170 ROIs with degree $k>208$, forming a core of density 0.8. In comparison, this is 56% denser than C4, which is the densest (0.51) of the network communities identified. The 100 leading hubs displayed in Fig.4a form a rich-club with an approximate density of 0.93; see their topographical location in Fig. 4c (right). As the threshold is increased, the number of hubs in the rich-club decreases. Mainly at the expense of cortical hubs, emphasizing the central role of subcortical hubs in the brain-wide network, Fig. 4c (lower panel) and Fig. 4d (right).
@@
The analysis performed here shows that, at the system?s level, the connectivity of the entire brain follows similar organization principles as those observed for the cortico-cortical connectivity alone: it is structured into a handful of network modules with the cross-modular connectivity centralized via hub ROIs, which form a spatially delocalized supramodule at the core, Fig. 4e. However, in this global case, we found that it is the subcortical ROIs—instead of cortical areas—which dominate the communication core of the network, while the cerebellum is excluded from the rich-club, Fig. 4e (left panel). Particularly relevant is the fact that no single subcortical structure centralizes the global pathways, but instead, subcortical hubs appear distributed across the different structures. This is further highlighted from the viewpoint of the network-based division, with the rich-club hubs evenly dispersed among the network communities.
@@
Figure 5: a, Schematic illustration of the integration capacity metric for a set of ROIs, in this case, the network hubs. Illustration for both the anatomical or the network divisions. Integration is defined as the joint response of a group of ROIs (the hubs), to stimuli applied in all other ROIs. b, Integration capacity of the 30 leading hubs, chosen from the global ranking (black bars) or chosen from the local ranking at each community (colored bars). Top, for the anatomical division into cortex, subcortex, and cerebellum; bottom, for the leading hubs within the network communities C1 – C5. Dashed lines indicate the average integration capacity of 30 randomly chosen ROIs, regardless of their hubness (average of 100 realizations). c, Integration capacity of the $n$ leading hubs, with $n$ from 1 to 170. Black lines, choosing hubs from their global ranking and, colored lines, choosing the hubs of a given community, see Fig. 4a (top and bottom). Notice that the sizes of cerebellum, C3, C4 and C5 are smaller than 170, thus, their last points reflect the integration capacity of the entire community. d, Schematic representation of the modular segregation metric, in both the anatomical and the network divisions. Segregation is defined as the conditional loss of communication between modules, given that a set of $n$ ROIs has been lesioned. In this case, a set of hubs is lesioned. e, Loss of communication between the modules after the leading 30 global hubs have been lesioned. Numbers indicate the loss between each pair of modules. f, Segregation suffered by the network after lesion of the 30 leading hubs, selected from the global ranking (black bars) or locally selected within a community (colored bars). Dashed lines indicate the average modular segregation after 30 randomly chosen ROIs are lesioned (average of 100 realizations). g, Modular segregation suffered by the network after the n leading hubs have been lesioned, with $n$ from 1 to 170. Black lines, choosing hubs from their global ranking and, colored lines, choosing them within each community. Notice that the sizes of cerebellum, C3, C4 and C5 are smaller than 170, thus, their last points reflect the segregation of the network after the entire community has been lesioned.
@@
We conclude by exploring the functional implications of the brain-wide architecture revealed so far and depicted in Fig. 4e. For that, we examine its role in balancing integration and segregation, and its vulnerability to targeted lesions. We evaluate integration and segregation following an in-silico perturbative approach based on network responses to external stimuli Zamora_Hubs_2010 . This approach accounts for the influence of brain regions on each other through all possible pathways, beyond one-to-one connections Gilson_DynCom_2019 ; Zamora_Chaos_2024 .
@@
Integration is defined as the joint response of a group of hubs to stimuli applied at the rest of the network, see Fig. 5a for an illustration. For example, taking the thirty leading hubs, either globally or locally per module, the integration capacity of subcortical hubs is clearly larger than that of cortical and cerebellar hubs, Fig. 5b (top). This uneven integration capacity is maintained when evaluated for groups of hubs of increasing size, by sequentially adding one hub at a time, Fig. 5c (top). The integration capacity of subcortical hubs (orange curve) is the largest across all group sizes—practically explaining the integration capacity of the global hubs (black curve), followed by that of cortical hubs (green curve). The integration capacity of the cerebellar hubs (blue) is very small in all cases. From the viewpoint of the network division, the integration capacity of hubs—selected by community—is more evenly distributed, Figs. 5b,c (bottom). Although C1 displays a larger integration than the rest and integration of C5 is clearly smaller. The integration of local hubs of the C1 – C4 communities remain close to each other regardless of the number of hubs taken, Fig. 5c (bottom).
@@
We define segregation as the loss of influence between communities, after a number of hubs has been lesioned Zamora_Hubs_2010 , see Fig. 5d for illustration. For quantification, the influence of one community on another is first measured in the original (healthy) network. Then, a number of hubs is lesioned and the cross-modular influences are measured again to evaluate the relative difference, from healthy to lesioned. When the thirty leading global hubs are lesioned, the segregation suffered by the network is drastic: 48–68% of communication is lost between the three anatomical communities and 38–63% among the network-based ones, Fig. 5e. The largest loss occurs between cortex and cerebellum (68%), given the circumstance that subcortical pathways are crucial for cortico-cerebellar communication. On the same order (63%) is the loss of communication incurred between C1 and C2 network communities. The total segregation after lesion of the 30 leading global hubs is of 50%, Fig. 5f (black bars). This level of segregation is striking taking into account that thirty hubs represent only 4% of all the ROIs. A comparable lesion of thirty ROIs chosen at random would only imply a segregation of 13% for either the anatomical or the network communities (dashed lines). Selecting the hubs per community, the largest segregation (50%) occurs when subcortical hubs are lesioned (top panel, orange bar), which is twice the segregation incurred by the lesion of cortical hubs (24%, green bar). Selecting hubs per network community, Fig. 5f (bottom), again we identify a more even segregation for targeted C1 – C4 hub lesions, as for their integration capacity.
@@ Discussion
The division of the brain into cortex, subcortical structures and cerebellum is an essential notion of the brain?s organization, rooted in their different cytoarchitecture, cellular composition and evolution. However, there is more to brain?s organization than meets the eye, hidden in the complex manner that white matter shapes the pathways between distant regions, and only discernible through suitable network analyses. Despite the tight functional interdependencies between cortex, subcortex and cerebellum, to date, a detailed description of the brain-wide communication network they form remains undisclosed. We employed high-quality tractography data from the Human Connectome Project to construct a structural connectome of the entire human brain (at the level of interconnected regions of interest) as a first step towards uncovering the principles of how cortex, subcortex and cerebellum are integrated.
First, we found that the brain is organized into network modules of mixed cortical, subcortical and/or cerebellar regions (Figs. 1 and 2). This mixture may facilitate the brain to collectively orchestrate a spectrum of functions ranging from basic motor control up to complex higher-order cognition. Second, we identified that the communication pathways between these modules are centralised through a number of highly connected ROIs forming a rich-club. Importantly, this rich-club is dominated by subcortical regions: 60% of the leading hubs are subcortical (Fig. 4) and their in-silico lesions greatly disrupt the communication throughout the network, significantly more than cortical lesions alone (Fig. 5e,f). Last, but not least, we found that subcortical structures are internally heterogeneous, with their ROIs displaying diverse local and global connectivity profiles and thus playing different roles in the network architecture (Fig. 3e). More so, the subcortical hubs are spatially segregated, meaning that all subcortical structures engage in the rich-club instead of one structure (e.g., thalamus or basal ganglia) centralising the communication pathways. Altogether, our results complement the traditional view of the large-scale organization of the human brain and invite to question the prevalent cortico-centric notion that regards subcortical structures and cerebellum as ‘mere’ supporters of cortical function. Instead, our results reveal a network architecture centered at the subcortex but made of transversal pathways.
@@
Subcortical regions are known to heavily interlink with the cortex via a multitude of pathways, for instance, the amygdala Phelps_Amygdala_2005 , the hippocampus Maller_Hippocampal_2019 ; Huang_Extensive_2021 , the basal ganglia Alexander_Basal_1986 ; Haber_Basal_2009 , the superior colliculus Benavidez_Superior_2021 , and the thalamus castro1997thalamocortical ; jones2001thalamic ; Haber_Basal_2009 . Also, the cerebellum is linked to the cortex via the cortico-ponto-cerebellar and the cerebello-thalamo-cortial pathways Ramnani_CorticoCereb_2006 ; Palesi_Pathways_2015 ; Palesi_Contralateral_2017 ; DAngelo_Physiology_2018 . The network communities here identified robustly underscore this mixed interconnectivity, as they show larger internal connection densities than separately the cortex, subcortex, and cerebellum but range over their boundaries, Fig. 1a,b.
The five communities identified contain ROIs associated with low-level sensory and motor processing, emotion, and high-level cognition including memory, attention, executive control, language and social cognition. Superimposed upon this extensive functional diversity, each module exhibits functional biases. Community C1 largely comprises of cerebellar and subcortical structures (mainly Diencephalon, brainstem and Pallidum), prominently featuring ROIs related to visceral or autonomic signal processing, emotion and motor control. C2 comprises primarily of anterior and posterior cortical regions together with subcortical structures (mostly thalamus, hippocampus, accumbens, caudate and amygdala), featuring memory-related ROIs. C3 and C4 are cortically dominated and show hemispheric lateralization. In line with prior reports on functional localization, the left-hemispheric community (C4) comprises more ROIs associated with language knecht2000language ; olulade2020neural and attention bartolomeo2019hemispheric than the right-hemispheric community (C3) whereas the right-hemispheric community comprises more ROIs associated with social cognition kobayakawa2018left ; gupta2025relationship . However, C5 is primarily made of medial cortical areas; it exhibits the largest proportion of ROIs related to social cognition and to internally oriented cognition.
@@
The existence and function of multimodal (transmodal) “convergence zones” has long been debated given the need of the brain to integrate “somewhere” sensory information from different modalities, initially processed by specialised regions Baars_Book_1988 ; Damasio_BrainBinds_1989 ; Tononi_Complexity_1994 ; mesulam1998sensation ; fuster2005cortex . Early structural connectivity data showed the presence of highly connected nodes (network hubs) in the neural system of the C. Elegans White_Elegans_1986 ; arenas2007elegans ; Varshney_Elegans_2011 and in the brains of macaques, cats and humans based primarily on cortico-cortical connectivity Scannell1993 ; Scannell1995 ; Sporns_Hubs_2007 ; Hagmann_Core_2008 ; Zamora_Chaos_2024 . It was soon identified that these network hubs were not only largely connected throughout sensory modalities, but were also densely interconnected with each other forming a higher hierarchical supramodule: a ‘rich-club’. A generalised assumption was drawn that such a rich-club may provide the structural backbone for centralised but distributed processing Zamora_Hubs_2010 ; Heuvel_HubsHuman_2011 previously postulated, for example, by the Global Workspace Theory Baars_GWT_2005 ; Shanahan_BroadcastNet_2010 .
@@
The importance of subcortical hubs has been recognised by studies focusing on specific structures, e.g. see Refs. [Bell_Subcortical_2016, ; Hansen_Brainstem_2024, ; cambareri2024subcortical, ]. Our results extend these reports by finding a rich-club core at the whole-brain scale which, crucially, is dominated by subcortical regions (Fig. 4a,c,d). Following our in-silico analyses of the propagation of stimuli, subcortical hubs emerged as the most integrative nodes within the global connectome (Fig. 5b,c) and their lesions led to significantly larger disruption of network communication than lesions of cortical or cerebellar hubs (Fig. 5f,g). These analyses thus contribute to a better understanding of clinical evidence that subcortical lesions are more likely to cause coma than cortical lesions Parvizi_Corticocentric_2009 ; posner2007plum ; fischer2016human , and that patients with subcortical lesions are less likely to recover consciousness than patients with cortical damage posner2007plum ; lutkenhoff2015thalamic ; annen2018regional .
The central role of the subcortical hubs for the brain-wide communication is remarkable considering that the subcortex only houses 0.8% of the neurons in the brain and 8% of the total brain mass VanEssen_Development_2018 . This disparity highlights the subcortical component of the global rich-club as a critical network bottleneck where widespread neural signals converge onto a rather small number of neurons whose activity is then broadcasted globally. Moreover, given the observed heterogeneity in network position among subcortical ROIs. Specifically, while individual subcortical structures harbor ROIs that serve as critical hubs within the brain-wide network, these same structures also contain ROIs exhibiting markedly lower connectivity and restricted inter-community integration. These observations raise compelling questions about the specific computations the subcortical structures and rich-club may perform. This intra-structural coexistence of highly integrative hubs alongside more segregated subregions may reflect a functional specialization, potentially delineating roles in localized information processing alongside broader network communication. Collectively, these observations underscore the imperative of considering the global connectome, and the topological position of brain regions within it, to achieve a more comprehensive understanding of their functional contributions.
@@
The construction of structural connectomes based on diffusion imaging are inherently constrained by the spatial resolution and technical capabilities of the equipment and methodology. The occurrence of false-negative and false-positive connections is a common issue when estimating structural connectivity from diffusion imaging. To mitigate this, the connectivity matrices were thresholded to preserve the original distribution of connection lengths, thus maintaining long-range fibers. Also, based on established anatomical knowledge direct cortico-cerebellar tracts, interhemispheric cerebellar connections, and ipsilateral cerebellar connections to the thalamus and brainstem were excluded. Some recent studies have investigated the structural connectivity between cortex and subcortex with a notable level of detail Benavidez_Superior_2021 ; alves2022subcortical ; giacometti2024differential ; Kumar_Thalamic_2023 . However, these works focused on selected subcortical structures. While this choice allowed high specificity of connections, they lack of the systems’ level viewpoint that we aimed at the present study, e.g., missing on the internal heterogeneity of connection profiles within these structure or the complex composition of the rich-club.
Obtaining both a high-level of detail and an overarching structural connectome is currently a major challenge for diffusion imaging. But the onset of ultra-high field imaging?scanning ex-vivo at high Teslas?will soon allow the description of the entire human structural connectivity at an unprecedented level of resolution beaujoin2018post ; lechanoine2021wikibrainstem . Together with the recent release of the whole-brain wiring map of fruit fly flywire2024whole ; lin2024network , we are about to undergo a breathtaking time to not only explore the brain?s organization, but also to inquire about its evolutionary mechanisms via comparative connectomics Suarez_Taxonomy_2022 ; Puxeddu_Relation_2024 . Despite the methodological precautions of this study, our findings represent an initial map and first wiring principles towards a comprehensive description of the brain?s overall architecture.

### v2

@@ Results
Figure 4: Hierarchical centralization of cross-modular connectivity. a, Analysis of the 100 lead hub ROIs of the brain-wide network. Bar panels display their degree—sorted in decreasing order—with the bars colored according to the community to which the ROI belongs, in the three different divisions. Pie charts show the fraction of leading hubs among the subsequent communities. b, Evidence for a rich-club supra-organization among the leading hubs. Left panel, conceptual illustration of a rich-club, formed when the leading hubs are densely connected with each other. Right panel, evolution of the link density for the subnetworks formed by ROIs with degree larger than $k$ (nodes with degree smaller than $k$ are removed), for all possible thresholds $k=0,1,\ldots,718$. Red curve is the evolution for the brain-wide network, gray curve is for degree-conserving randomized samples (100 independent realizations each point). The threshold for $k=0$ represents the entire network, thus, the curves begin from a density of 0.2. Their monotonic increase up to a density of 1.0 evidences that the leading hubs eventually form an all-to-all connected core, centralizing the communication pathways. c, Topographical location of the rich-club hubs and their mutual links, for two different density thresholds: at $k>256$ (subgraph density of 0.93) and $k>300$ (subgraph density of 0.99). Cortical ROIs are shown as green dots and subcortical ROIs in orange. d, The composition of the rich-club at different levels of internal density, from 0.8 to 1.0 (degree thresholds, $k\geq 208$ to $k=365$), displayed for the anatomical (left or right cortex, subcortex, cerebellum) and the network divisions (C1 – C5). Bars indicate the number of ROI hubs forming the rich-club at each density level, and how many of them belong to the different anatomical or network communities. e, Schematic representation of the modular and hierarchical organization of the brain-wide structural connectivity, in the light of either the anatomical or the optimal network divisions.
@@
The results of Fig. 3 highlight the presence of hub ROIs in the cortex and all subcortical structures. We now examine how these hubs help elucidate the brain-wide network architecture. Ranking the 100 most connected ROIs together, Fig. 4a top panel, reveals that the leading hubs are mainly subcortical (60%), followed by cortical hubs (40%). Remarkably, no cerebellar ROI is present among the top global hubs. Distinguishing these leading hubs by subcortical structure (middle panel) confirms that all subcortical structures contain hub ROIs (regardless of their size) and thus, no single structure can be considered responsible for the centralization of the brain-wide connectivity. From the network-based division (bottom panel) the leading hubs are broadly distributed with the C1 – C4 communities containing 19% – 29% of the hubs each.
@@
Additionally, we find clear evidence that these leading hubs form a rich-club supra-community, see Fig. 4b (left panel) for an illustration. By iteratively removing the ROIs with less connections, the internal density of the remaining subnetwork (made of the ROIs with larger degree) rapidly increases, Fig. 4b (right panel). There are 170 ROIs with degree $k>208$, forming a core of density 0.8. In comparison, this is 56% denser than C4, which is the densest (0.51) of the five network communities. The 100 leading hubs displayed in Fig. 4a form a rich-club with internal density of 0.93; their topographical location is shown in Fig. 4c.
As the threshold is increased, the number of hubs in the rich-club decreases, Fig. 4d. This happens mainly at the expense of cortical hubs (left panel), emphasizing the central role of the subcortex in the brain-wide connectome. However, the representation of network communities in the rich-club remains balanced across thresholds (right panel).
@@
The analysis performed here shows that, at the systems level, the connectivity of the entire brain follows similar organization principles as those observed for the cortico-cortical connectivity alone Zamora-López et al. (2010, 2011); van den Heuvel and Sporns (2011): it is structured into a handful of network modules with the cross-modular connectivity centralized via hub ROIs, which form a spatially delocalized supramodule at the core, Fig. 4e. However, in this global case, we found that it is the subcortical ROIs—instead of the cortical areas—which dominate the communication core of the network, while the cerebellum is excluded from the rich-club, Fig. 4e (left panel). Particularly relevant is the fact that no single subcortical structure centralizes the global pathways, but instead, subcortical hubs appear distributed across the different structures. This is further highlighted from the viewpoint of the network-based division, with the rich-club hubs evenly dispersed among the network communities.
@@
Figure 5: Integrative and segregative consequences of the brain-wide network architecture. a, Schematic illustration of the integration capacity metric for a set of ROIs, in this case, the network hubs. Illustration for both the anatomical or the network divisions. Integration is defined as the joint response of a group of ROIs (the hubs), to stimuli applied in all other ROIs. b, Integration capacity of the 30 leading hubs, chosen from the global ranking (black bars) or chosen from the local ranking at each community (colored bars). Top, for the anatomical division into cortex, subcortex, and cerebellum; bottom, for the leading hubs within the network communities C1 – C5. Dashed lines indicate the average integration capacity of 30 randomly chosen ROIs, regardless of their hubness (average of 100 realizations). c, Integration capacity of the $n$ leading hubs, with $n$ from 1 to 170. Black lines, choosing hubs from their global ranking and, colored lines, choosing the hubs of a given community, see Fig. 4a (top and bottom). Notice that the sizes of cerebellum, C3, C4 and C5 are smaller than 170, thus, their last points reflect the integration capacity of the entire community. d, Schematic representation of the modular segregation metric, in both the anatomical and the network divisions. Segregation is defined as the conditional loss of communication between modules, given that a set of $n$ ROIs has been lesioned. In this case, a set of hubs is lesioned. e, Loss of communication between the modules after the leading 30 global hubs have been lesioned. Numbers indicate the loss between each pair of modules. f, Segregation suffered by the network after lesion of the 30 leading hubs, selected from the global ranking (black bars) or locally selected within a community (colored bars). Dashed lines indicate the average modular segregation after 30 randomly chosen ROIs are lesioned (average of 100 realizations). g, Modular segregation suffered by the network after the n leading hubs have been lesioned, with $n$ from 1 to 170. Black lines, choosing hubs from their global ranking and, colored lines, choosing them within each community. Notice that the sizes of cerebellum, C3, C4 and C5 are smaller than 170, thus, their last points reflect the segregation of the network after the entire community has been lesioned.
@@
We conclude this study by exploring functional implications of the brain-wide architecture uncovered and depicted in Fig. 4e. For that, we examine its capacity to host integrative and segregative processes, and its vulnerability to targeted lesions. We evaluate integration and segregation following an in-silico perturbative approach that assesses network responses to external stimuli Zamora-López et al. (2010). This approach accounts for how brain regions influence each other through all possible pathways, beyond one-to-one interactions Gilson et al. (2019); Zamora-López and Gilson (2024).
@@
Integration is defined as the joint response of a set of hubs to stimuli applied at the rest of the network, see Fig. 5a for an illustration. Considering the thirty leading hubs, either globally or locally per module, the integration capacity of subcortical hubs is clearly larger than that of cortical and cerebellar hubs, Fig. 5b (top). We estimate the integration capacity for different hub set sizes by sequentially adding one hub at a time, Fig. 5c. The integration capacity of subcortical hubs (top panel, orange curve) is the largest across all hub set sizes, practically explaining the integration capacity of the global hubs (black curve), followed by that of cortical hubs (green curve). The integration capacity of the cerebellar hubs (blue) is very small in all cases. From the viewpoint of the network division, the integration capacity of hubs (selected by community) is more evenly distributed, Figs. 5b,c (bottom panels). Although C1 displays larger integration than the rest, the curves for integration capacity of C1 – C4 communities remain close to each other at all hub set sizes considered, Fig. 5c (bottom).
@@
We define segregation as the loss of influence between communities, after a number of hubs has been lesioned Zamora-López et al. (2010), see Fig. 5d for illustration. For example, the influence of C1 on C2 is first measured in the original (healthy) network. Then, a set of hubs is lesioned and the influence is re-evaluated. Segregation is determined as the relative difference (loss) of influence between C1 and C2, from healthy to lesioned. When the thirty leading global hubs are lesioned, the segregation suffered by the network is drastic: 48 – 68% of influence is lost between the three anatomical communities and 38 – 63% among the network-based ones, Fig. 5e. The largest loss occurs between cortex and cerebellum (68%). On the same order (63%) is the loss of communication incurred between C1 and C2 network communities. The total segregation incurred after lesion of the 30 leading global hubs is 50%, Fig. 5f (black bars). This level of segregation is striking taking into account that thirty hubs represent only 4% of all the ROIs. Selecting the hubs per community, the largest segregation (50%) occurs when subcortical hubs are lesioned (top panel, orange bar), which is twice the segregation incurred by the lesion of cortical hubs (24%, green bar). Selecting hubs following network communities, Fig. 5f (bottom), again we identify a more even segregation after targeted C1 – C4 hub lesions, as for their integration capacity.
@@ Discussion
A major challenge for neuroscience today is to better understand how cortex, subcortex and cerebellum work together. This anatomical division is an essential notion of the brain’s organization, evidenced by their differentiated cytoarchitecture, cellular composition and evolution Pinel and Barnes (2018). However, there is more to the brain’s structure than meets the eye, hidden in the complex manner that white matter shapes the pathways between distant regions, and only discernible through suitable network analyses. Here, we employed high-quality diffusion imaging and tractography from the Human Connectome Project, and a parcellation in 718 regions of interest (ROIs) Ji et al. (2019), in order to uncover the macroscopic connectional architecture of the entire brain.
@@
Subcortical regions are known to heavily interlink with the cortex via a multitude of pathways involving the amygdala Phelps and LeDoux (2005), the hippocampus Maller et al. (2019); Huang et al. (2021), the basal ganglia Alexander et al. (1986); Haber and Calzavara (2009), the superior colliculus Benavidez et al. (2021), and the thalamus Castro-alamancos and Connors (1997); Jones (2001); Haber and Calzavara (2009). On the other hand, cerebellum and cortex are indirectly connected through the subcortex: via the cerebello-thalamo-cortical and the cortico-ponto-cerebellar pathways Palesi et al. (2017). Here, we found the brain to be organized into five macroscopic network modules of mixed ROIs, underlining that those extensive pathways form ordered brain-wide circuits surpassing the macro-anatomical boundaries between cortex, subcortex, and cerebellum (see Figs. 1a,b and 2).
The five network modules contain ROIs associated with low-level sensory and motor processing, emotion, and high-level cognition including memory, attention, executive control, language and social cognition. Superimposed upon this extensive functional diversity, each module exhibits a degree of functional bias. Community C1 largely comprises cerebellar and subcortical structures (diencephalon, brainstem and pallidum), prominently featuring regions related to visceral or autonomic signal processing, emotion and motor control. C2 comprises primarily anterior and posterior cortical regions together with subcortical structures (mostly thalamus, hippocampus, accumbens, caudate and amygdala), featuring memory-related regions. C3 and C4 are cortically dominated and show hemispheric lateralization. In line with prior reports on functional localization, the left-hemispheric community (C4) comprises more ROIs associated with language Knecht et al. (2000); Olulade et al. (2020) and attention Bartolomeo and Malkinson (2019) than the right-hemispheric community (C3), whereas the latter comprises more regions associated with social cognition Kobayakawa (2018); Gupta et al. (2025). However, C5—primarily composed of medial cortical areas—exhibits the largest proportion of ROIs related to social cognition and to internally oriented cognition.
@@
Given the need of the brain to integrate sensory information from different modalities, initially processed by specialized regions, the existence and function of multimodal “convergence zones” has long been debated Baars (1988); Damasio (1989); Tononi et al. (1994); Mesulam (1998); Fuster (2005). Early studies showed the presence of hubs in the cortico-cortical connectomes of macaques, cats and humans Scannell and Young (1993); Scannell et al. (1995); Sporns et al. (2007); Hagmann et al. (2008); Zamora-López et al. (2010), which are broadly connected across sensory modalities, and densely interconnected with each other, forming a hierarchical supra-module (a ‘rich-club’). Then, the conjecture was put forward that such a rich-club may provide the structural backbone for centralized but distributed processing Zamora-López et al. (2010); van den Heuvel and Sporns (2011) as postulated, for example, by the Global Workspace Theory Baars (2005); Shanahan (2010).
@@
The presence of subcortical hubs has been reported by studies focusing on specific structures Bell and Shine (2016); Hansen et al. (2024); Cambareri et al. (2024). Our results complete these reports by identifying a rich-club at the core of the brain’s global pathways which, crucially, is dominated by subcortical regions (Fig. 4a,c,d). Our in-silico analyses, based on the propagation of external stimuli throughout the brain-wide connectome, showed that subcortical hubs emerged as the most integrative nodes (Fig. 5b,c) and their lesions led to significantly larger disruption of network communication than lesions of cortical or cerebellar regions (Fig. 5f,g). These results are consistent with clinical evidence that patients with subcortical lesions are more likely to suffer from coma Parvizi (2009); Posner et al. (2007); Fischer et al. (2016), and are less likely to recover consciousness than patients with cortical damage alone Posner et al. (2007); Lutkenhoff et al. (2015); Annen et al. (2018).
The central role of the subcortical hubs is remarkable considering that the subcortex only houses 0.8% of the neurons in the brain van Essen et al. (2018) while the cerebellum, containing 80% of the neurons, did not display any global hub. This disparity highlights the subcortical component of the global rich-club as a critical network bottleneck, where widespread neural signals converge onto a rather small number of neurons, whose activity is then broadcasted globally.
A surprising observation from our analyses is that, not only is the rich-club subcortically dominated but all subcortical structures partly contribute to it (Fig. 4a, middle panel). Common understanding of brain pathways used to assign, for example, a central role to the thalamus because it is the entry point for all sensory inputs to the cortex, while other subcortical structures were usually thought to deal with more specialized functions. Instead, here we have shown that the center of the brain-wide network is shared by all subcortical structures. This is possible because subcortical structures contain ROIs of diverse connectivity profiles (Fig. 3e) implying that those ROIs serve different roles for the global network architecture. Within each subcortical structure we find ROIs of markedly lower (local) connectivity, likely serving more specialized functions, while other ROIs are broadly connected across the network modules and take part in the global rich-club. This internal diversity within subcortical structures raises compelling questions about their level of specialization. The presence of both locally and globally connected ROIs opens the door to clarify how individual subcortical structures may become involved in various functions, from perception to emotional and cognitive processes.
@@
The construction of structural connectomes based on diffusion imaging is inherently constrained by the spatial resolution and technical capabilities of the imaging procedure and subsequent preprocessing. The occurrence of false-negative and false-positive connections is a common issue when estimating structural connectivity from diffusion imaging. To mitigate this, the connectivity matrices were thresholded while preserving the original distribution of connection lengths, thus maintaining long-range fibers. Also, based on established anatomical knowledge, direct cortico-cerebellar tracts, inter-hemispheric cerebellar connections, and ipsilateral cerebellar connections to the thalamus and brainstem were excluded. Some recent studies have investigated the structural connectivity between cortex and subcortex with a notable level of detail Benavidez et al. (2021); Alves et al. (2022); Giacometti et al. (2024); Kumar et al. (2023). However, these works focused on selected subcortical structures. While this choice allowed high specificity of connections, they lack the systems level viewpoint of the present study, e.g., missing the internal heterogeneity of connection profiles within these structures, or the complex composition of the rich-club.
Obtaining structural connectomes with both a high-level of detail and an overarching scale is currently a major challenge for diffusion imaging. But the onset of ultra-high field imaging—scanning ex-vivo—will soon allow the description of the human structural connectivity at an unprecedented level of resolution and completeness Plantinga et al. (2016); Roebroeck et al. (2018); Beaujoin et al. (2018); Lechanoine et al. (2021). Together with the recent release of the whole-brain wiring map of the fruit fly Consortium et al. (2024); Lin et al. (2024), we are about to undergo a breathtaking time to not only explore the brain’s organization, but also to inquire about its evolutionary mechanisms via comparative connectomics Suarez et al. (2022); Puxeddu et al. (2024). Despite the methodological precautions of this study, our findings represent an initial map and first wiring principles towards a comprehensive description of the brain’s overall architecture.